OpenStudy (anonymous):

Explain me this: U²-(6/7)u+(1-7)+V²=0 Becomes, (U-(3/7))²+V²= (4/49)-(1/7) I can't figure out this...

OpenStudy (ranga):

It is called completing the square. For example if you have x^2 + bx, you can complete the square as follows: x^2 + bx = (x + b/2)^2 - (b/2)^2 You can expand the terms on the right and simplify to verify the right hand side is same as the left hand side.

OpenStudy (ranga):

Here you have U^2 - (6/7)U Take the coefficient of U and divide it by 2. That is the term that will go inside the square: -6/7 divided by 2 = -3/7 U^2 - (6/7)U = (U - 3/7)^2 - (-3/7)^2

OpenStudy (anonymous):

Thank you!!! :)

OpenStudy (ranga):

you are welcome.